# Vision-Only Gaussian Splatting for Collaborative Semantic Occupancy Prediction

### v1

@@ Introduction
Multi-agent collaborative perception (CP), also known as cooperative perception, significantly enhances transportation safety, mobility, and efficiency by enabling connected vehicles to integrate multiple viewpoints through Vehicle-to-Everything (V2X) communication, forming a comprehensive understanding of their environment (Huang et al. 2023a). Early CP studies primarily focused on established single-agent tasks such as 3D object detection (3DOD) (Liu et al. 2020; Hu et al. 2022; Ding et al. 2025; Zhang et al. 2024; Xu et al. 2022b, 2023) and 2D bird’s-eye-view (BEV) semantic segmentation (Wang et al. 2020; Li et al. 2021), typically by fusing latent BEV features (Figure 1 top). Historically reliant on LiDAR point clouds for their precise geometric measurements, recent advances in vision-only methods have shown competitive performance. For example, CoCa3D (Hu et al. 2023) demonstrates that camera-based 3DOD can match or surpass LiDAR-based approaches, highlighting cameras as viable, cost-effective sensors for complex 3D scene understanding.
@@
However, tasks like 3DOD and BEV segmentation simplify the perception of the scene by omitting crucial 3D semantic details, and BEV-based features inherently suffer from information loss due to height compression. This limitation hinders holistic scene understanding and downstream decision-making. 3D semantic occupancy prediction (SOP) (Roldao, De Charette, and Verroust-Blondet 2022; Tian et al. 2023; Li et al. 2024) bridges this gap by predicting the occupancy status of each voxel in the surrounding 3D space, thus delivering fine-grained semantic and geometric scene information. Vision-centric SOP methods (Cao and De Charette 2022; Song et al. 2017; Chen et al. 2020; Li et al. 2023; Wei et al. 2023; Huang et al. 2023b) have shown promising results on large-scale datasets, yet extending SOP to collaborative scenarios remains largely unexplored. CoHFF (Song et al. 2024) introduced the first collaborative SOP framework, demonstrating the benefits of V2X feature fusion. Nevertheless, it relies on planar tri-plane features (Huang et al. 2023b; Fridovich-Keil et al. 2023) (Figure 1 middle) which requires explicit depth supervision, and employs complex multi-stage, multi-task training pipelines, which introduces computational inefficiencies and complicates cross-agent feature alignment. Inspired by advances in 3D Gaussian splatting (Kerbl et al. 2023), sparse, object-centric representations have emerged as promising alternatives to traditional dense voxel or planar encodings for SOP (Li et al. 2023; Wei et al. 2023; Huang et al. 2023b). Driving scenes are highly sparse—most voxels are empty—so anisotropic Gaussians can shape and orient along semantic object surfaces, concentrating capacity where geometry and semantics lie while keeping free space compact, yielding an efficient 3D scene model. GaussianFormer (Huang et al. 2024b, a) represents scenes using sets of 3D Gaussians, each defined by a mean, covariance, and semantic label; these Gaussians are refined via deformable attention and rendered to a voxel grid through Gaussian-to-voxel splatting.
@@
In this work, we propose the first Vision-Only Gaussian Splatting for SOP framework that can be efficiently trained end-to-end in a single stage. Our method leverages sparse 3D semantic Gaussians as intermediate representations shared among vehicles, explicitly encoding both geometric and semantic information. By transmitting rigid transforms of Gaussians only within each ego vehicle’s region of interest (Figure 1 bottom), our approach effectively integrates multi-agent viewpoints, overcoming occlusions. Additionally, we introduce a novel cross-agent Gaussian fusion module that refines noisy, redundant, and conflicting Gaussians through neighborhood-based fusion, enabling accurate and realistic semantic occupancy prediction. To the best of our knowledge, we are the first to utilize 3D Gaussian splatting for multi-agent collaborative 3D semantic occupancy prediction. We summarize our contributions as:
@@
Semantic occupancy prediction (SOP), also known as semantic scene completion (SSC), estimates per-voxel occupancy together with semantic categories, yielding a volumetric representation of geometry and semantics. Since Tesla described an occupancy network for Full Self-Driving (Tesla 2022), research on 3D occupancy for autonomous driving has steadily accelerated. Methods differ by sensing modality: some use LiDAR point clouds (Zhao et al. 2025) or 4D radar tensors (4DRTs) (Ding et al. 2024), while many recent systems are vision-only and lift image features into 3D using learned geometric priors (Cao and De Charette 2022; Song et al. 2017; Chen et al. 2020; Li et al. 2020; Huang et al. 2023b, 2024b; Wei et al. 2023; Zhang, Zhu, and Du 2023; Li et al. 2023). Most prior work addresses single-vehicle perception. Extending SOP to multi-agent collaborative perception centers on two design decisions: the communication representation and the fusion strategy for processing aggregated information under bandwidth limits and field-of-view differences. Collaborative SOP is only beginning to be explored; CoHFF (Song et al. 2024) reports clear benefits from sharing but relies on multi-stage training and depth estimation, which increases system complexity and complicates cross-agent feature alignment.
@@
Collaborative perception for connected autonomous vehicles (CAVs) uses V2X communication and data fusion to improve scene understanding. Fusion is commonly grouped into early, intermediate, and late (Huang et al. 2023a). Early fusion shares raw sensor data (e.g., point clouds), which is bandwidth demanding and raises privacy concerns (He et al. 2021). Late fusion exchanges object lists, which is lightweight but discards fine details (Shi et al. 2022; Zhu et al. 2024). Most recent systems adopt intermediate fusion, which exchanges latent features and retains more task signal while avoiding raw data transfer (Wang et al. 2020; Hu et al. 2022; Xu et al. 2022a, b, 2024; Ding et al. 2025). While many works target LiDAR or hybrid setups for 3D detection or BEV segmentation (Xu et al. 2023; Ding et al. 2025; Zhang et al. 2024; Zimmer et al. 2024; Li et al. 2021; Hu et al. 2022; Yang et al. 2023), vision-only collaboration is increasingly practical for large-scale deployment. CoHFF brought collaborative semantic occupancy prediction into focus by demonstrating SOP benefits from V2X feature sharing and fusion (Song et al. 2024).
@@
3D Gaussian Splatting (3DGS) represents scenes with anisotropic Gaussians and renders them by splatting for novel view synthesis (Kerbl et al. 2023). Subsequent work adapts Gaussian splatting to perception tasks, including object detection (Cao, Jv, and Xu 2024; Yan, Zheng, and Duan 2024). For semantic occupancy, the GaussianFormer family splats Gaussians into voxels (Huang et al. 2024b, a). Prior work treats Gaussians as an internal single‑agent representation. We instead introduce Gaussians as the communication medium for collaboration. This choice provides compact messages, closed‑form rigid alignment across agents, region‑of‑interest culling, and explicit geometry that reduces reliance on separate depth supervision. Our method builds communication and cross-agent fusion for Gaussian primitives.
@@
$\mathbf{F}_{i}$ $=f_{\text{enc}}(\mathbf{O}_{i}),\enspace i\in\mathcal{S}\enspace,$ (1)
$\mathbf{F}_{j\rightarrow i}$ $=\Gamma_{j\rightarrow i}(\mathbf{F}_{j}),\enspace j\in\mathcal{S}\enspace,$ (2)
$\mathbf{H}_{i}$ $=f_{\text{fusion}}(\{\mathbf{F}_{j\rightarrow i}\}_{j\in\mathcal{S}})\enspace,$ (3)
$\mathbf{B}_{i}$ $=f_{\text{head}}(\mathbf{H}_{i})\enspace,$ (4)
@@
$\max_{\theta}\sum_{i}g\left(\Phi_{\theta}(\mathbf{O}_{i},\{\mathbf{F}_{j\rightarrow i}\}_{j\in\mathcal{S}}),\mathbf{\hat{B}}_{i}\right),\enspace\text{s.t.}\enspace\sum_{j}\left|\mathbf{F}_{j\rightarrow i}\right|\leq\beta\enspace,$ (5)
@@
Considering practical bandwidth limits, extending voxel-based single-agent SOP methods (e.g.,(Wei et al. 2023; Li et al. 2023)) to collaboration is not viable: transmitting dense voxel features in $\mathbb{R}^{X\times Y\times Z\times D}$ (where $D$ is the hidden dimension) is too costly. Drawn inspiration from TPVFormer (Huang et al. 2023b), CoHFF proposes to transmit features from orthogonal planes $\mathbf{P}^{xz}$ and $\mathbf{P}^{yz}$, with features $\mathbf{F}^{\mathbf{P}^{xz}}\in\mathbb{R}^{X\times Z\times F}$ and $\mathbf{F}^{\mathbf{P}^{yz}}\in\mathbb{R}^{Y\times Z\times F}$. This reduces the communication volume from $XYZD$ to $(XZ+YZ)D$ and makes collaborative SOP feasible.
@@
Inspired by GaussianFormer (Huang et al. 2024b), we opt to represent a scene with a set of 3D Gaussian primitives $\mathcal{G}=\{\mathbf{G}_{i}\in\mathbb{R}^{d}|i=1,\ldots,P\}$. Distinct from GaussianFormer, we make the set of 3D Gaussian primitives serve as a communication medium for collaborative perception. Specifically, each $\mathbf{G}_{i}$ describes a local region with its mean $\mathbf{m}_{i}\in\mathbb{R}^{3}$, scale $\mathbf{s}_{i}\in\mathbb{R}^{3}$, rotation $\mathbf{r}_{i}\in\mathbb{R}^{4}$, opacity $a_{i}\in\mathbb{R}^{1}$, and semantics $\mathbf{c}_{i}\in\mathbb{R}^{|\mathcal{C}|}$. These primitives are interpreted as local semantic Gaussian distributions which contribute to the overall occupancy prediction for any given 3D location $\mathbf{x}\in\mathbb{R}^{3}$ through additive aggregation:
$\hat{\mathbf{o}}(\mathbf{x};\mathcal{G})=\sum_{i=1}^{P}\mathbf{g}_{i}(\mathbf{x};\mathbf{m}_{i},\mathbf{s}_{i},\mathbf{r}_{i},a_{i},\mathbf{c}_{i})\enspace,$ (6)
@@
$\mathbf{g}(\mathbf{x};\mathbf{G})=a\cdot\exp\left(-\frac{1}{2}(\mathbf{x}-\mathbf{m})^{\text{T}}\mathbf{\Sigma}^{-1}(\mathbf{x}-\mathbf{m})\right)\mathbf{c}\enspace,$ (7)
$\mathbf{\Sigma}=\mathbf{RSS}^{\top}\mathbf{R}^{\top}\enspace,\enspace\mathbf{S}=\text{diag}(\mathbf{s})\enspace,\enspace\mathbf{R}=\text{q2r}(\mathbf{r})\enspace,$ (8)
@@
From a single-agent perspective, the pipeline first estimates 3D Gaussian primitives with an image-to-Gaussian module, then renders semantic occupancy via Gaussian-to-voxel splatting. We adopt both components from GaussianFormer (Huang et al. 2024b). We use 3D Gaussian primitives as a compact communication interface for collaborative perception. A Gaussian packaging module transforms each neighbor’s Gaussians into the ego frame and culls them to the ego region of interest, reducing the transmitted data. A cross-agent Gaussian fusion module then refines the ego set by aggregating consistent multi-view evidence and down-weighting noisy or low-opacity inputs. Collaborative semantic occupancy is obtained by splatting the refined ego Gaussians. The system is illustrated in Fig. 2.
Figure 2: Overview of the proposed pipeline. An initial set of randomly initialized 3D Gaussians is refined by an Image-to-Gaussian module (Huang et al. 2024b) that attends to multi-scale image features, producing single-agent Gaussians. Neighbor agents (top and bottom) are rigidly transformed into the ego frame (middle) and culled to the ego region of interest; the black dashed box marks the Gaussians that lie within the ego ROI and are packaged and transmitted to the ego. A cross-agent Gaussian fusion module aggregates these with the ego set. The fused Gaussians are then rendered to semantic occupancy via Gaussian-to-voxel splatting (Huang et al. 2024b). For clarity, the figure shows the zero-shot variant.
@@
Unlike dense voxel grid features (Wei et al. 2023; Li et al. 2023) or planar features (Song et al. 2024; Huang et al. 2023b), Gaussian primitives model ellipsoidal probability densities in $\mathbb{R}^{3}$, which remain well defined under rigid transforms (Kerbl et al. 2023; Huang et al. 2024b), and thus cross-agent 3D Gaussian alignment and transmission are straightforward and interpretable. Specifically, let the known extrinsic from agent $j$ to agent $i$ be $T_{ij}(x)=U_{ij}x+t_{ij}$ with $U_{ij}\in\mathrm{SO}(3)$, $t_{ij}\in\mathbb{R}^{3}$, and quaternion $q_{ij}$ such that $\text{q2r}(q_{ij})=U_{ij}$.
@@
$\mathbf{m}^{\prime}$ $=U_{ij}\,\mathbf{m}+t_{ij}\enspace,$ (9)
$\mathbf{\Sigma}^{\prime}$ $=U_{ij}\,\mathbf{\Sigma}\,U_{ij}^{\top}\;=\;\big{(}U_{ij}\mathbf{R}\big{)}\,\mathbf{S}\mathbf{S}^{\top}\,\big{(}U_{ij}\mathbf{R}\big{)}^{\top}\enspace.$ (10)
@@
$\mathbf{m}^{\prime}=U_{ij}\mathbf{m}+t_{ij}\enspace,\enspace\mathbf{s}^{\prime}=\mathbf{s}\enspace,\enspace$
$\mathbf{r}^{\prime}=q_{ij}\otimes\mathbf{r}\enspace,\enspace a^{\prime}=a\enspace,\enspace\mathbf{c}^{\prime}=\mathbf{c}\enspace,$ (11)
@@
$\mathcal{G}_{j\rightarrow i}\;=\;\{\,\mathbf{G}\in\mathcal{G}_{j}\;|\;\mathbf{m}^{\prime}\in\mathbf{ROI}_{i}\,\}\enspace.$ (12)
@@
$\mathcal{G}^{\mathrm{stack}}_{i}\;=\;\mathcal{G}_{i}\;\cup\;\{\mathcal{G}_{j\rightarrow i}\}_{j\in\mathcal{S}}\enspace,$ (13)
@@
Unlike the refinement step in the image-to-Gaussian module of GaussianFormer (Huang et al. 2024b), our fusion module decodes neighborhood-conditioned proposals, pools them across nearby Gaussians to suppress noise and enforce consistency, and then blends the pooled update with the ego Gaussian attributes.
@@
Table 1: Comparison for semantic occupancy prediction. In the collaborative setting, our naive fusion and learned fusion variants achieve better IoU and mIoU than CoHFF (Song et al. 2024). The best is in bold and the second best is underlined.
@@
$\mathcal{G}_{\text{nbgr},k}=\{\,\mathbf{G}_{j}\in\mathcal{G}_{\text{nbgr}}\mid\|\mathbf{m}_{j}-\mathbf{m}_{k}\|_{2}\leq\rho\,\}\enspace.$ (14)
@@
$\mathbf{f}^{\mathrm{ego}}_{k}$ $=\big{[}\,\mathbf{m}_{k}^{\top},\ \mathbf{s}_{k}^{\top},\ \mathbf{r}_{k}^{\top},\ a_{k},\ \mathbf{c}_{k}^{\top}\,\big{]}\enspace,$ (15)
$\mathbf{f}^{\mathrm{rel}}_{k,j}$ $=\big{[}\,(\mathbf{m}_{j}{-}\mathbf{m}_{k})^{\top},\ (\mathbf{s}_{j}{-}\mathbf{s}_{k})^{\top},\ |\langle\mathbf{r}_{j},\mathbf{r}_{k}\rangle|,\ a_{j},\ \mathbf{c}_{j}^{\top}\,\big{]}\enspace,$
$\mathbf{z}_{k,j}$ $=\big{[}\,\mathbf{f}^{\mathrm{ego}}_{k}\ \big{|}\ \mathbf{f}^{\mathrm{rel}}_{k,j}\,\big{]}\enspace,$
@@
$\mathbf{u}_{k,j}\coloneq\big{(}\Delta\mathbf{m}_{k,j},\ \mathbf{s}^{\star}_{k,j},\ \mathbf{r}^{\star}_{k,j},\ a^{\star}_{k,j},\ \mathbf{c}^{\star}_{k,j}\big{)}\;=\;\text{MLP}(\mathbf{z}_{k,j})\enspace,$ (16)
@@
$w_{k,j}=\mathrm{softmax}_{j}\!\left(\frac{\langle Q\,\mathbf{x}^{\text{ego}}_{k},\ K\,\mathbf{x}^{\text{rel}}_{k,j}\rangle}{\sqrt{d}}\right)\enspace,$ (17)
@@
$\bar{\mathbf{u}}_{k}=\sum_{G_{j}\in\mathcal{G}_{\text{nbgr},k}}w_{k,j}\,\mathbf{u}_{k,j}\enspace.$ (18)
@@
$\hat{\mathbf{m}_{k}}$ $=\mathbf{m}_{k}+\overline{\Delta\mathbf{m}}_{k}\enspace,\qquad\hat{\mathbf{s}_{k}}\;=\;\overline{\mathbf{s}^{\star}}_{k}\enspace,$ (19)
$\hat{\mathbf{r}_{k}}$ $=\overline{\mathbf{r}^{\star}}_{k}\enspace,\qquad\hat{a}_{k}\;=\;\overline{a^{\star}}_{k}\enspace,$
$\hat{\mathbf{c}_{k}}$ $=\alpha_{k}\,\mathbf{c}_{k}+(1{-}\alpha_{k})\,\overline{\mathbf{c}^{\star}}_{k}\enspace,$
@@
$\alpha_{k}=\frac{\mathrm{conf}(\mathbf{c}_{k})}{\mathrm{conf}(\mathbf{c}_{k})+\mathrm{conf}(\overline{\mathbf{c}^{\star}}_{k})}\,\enspace,\quad\mathrm{conf}(v)=\max\!\Big{(}\frac{v}{\mathbf{1}^{\top}v+\varepsilon}\Big{)}\enspace.$ (20)

### v2

@@ Introduction
Multi-agent collaborative perception (CP), also known as cooperative perception, significantly enhances transportation safety, mobility, and efficiency by enabling connected vehicles to integrate multiple viewpoints through Vehicle-to-Everything (V2X) communication, forming a comprehensive understanding of their environment (cp_survey). Early CP studies primarily focused on established single-agent tasks such as 3D object detection (3DOD) (who2com; where2comm; cppc; ermvp; v2xvit; v2v4real) and 2D bird’s-eye-view (BEV) semantic segmentation (v2vnet; disconet), typically by fusing latent BEV features (Figure 1 top). Historically reliant on LiDAR point clouds for their precise geometric measurements, recent advances in vision-only methods have shown competitive performance. For example, CoCa3D (coca3d) demonstrates that camera-based 3DOD can match or surpass LiDAR-based approaches, highlighting cameras as viable, cost-effective sensors for complex 3D scene understanding.
@@
However, tasks like 3DOD and BEV segmentation simplify the perception of the scene by omitting crucial 3D semantic details, and BEV-based features inherently suffer from information loss due to height compression. This limitation hinders holistic scene understanding and downstream decision-making. 3D semantic occupancy prediction (SOP) (ssc-survey; occ3d; sscbench) bridges this gap by predicting the occupancy status of each voxel in the surrounding 3D space, thus delivering fine-grained semantic and geometric scene information. Vision-centric SOP methods (monoscene; ssc1; ssc2; voxformer; surroundocc; tpvformer) have shown promising results on large-scale datasets, yet extending SOP to collaborative scenarios remains largely unexplored. CoHFF (cohff) introduced the first collaborative SOP framework, demonstrating the benefits of V2X feature fusion. Nevertheless, it relies on planar tri-plane features (tpvformer; kplanes) (Figure 1 middle) which requires explicit depth supervision, and employs complex multi-stage, multi-task training pipelines, which introduces computational inefficiencies and complicates cross-agent feature alignment. Inspired by advances in 3D Gaussian splatting (3dgs), sparse, object-centric representations have emerged as promising alternatives to traditional dense voxel or planar encodings for SOP (voxformer; surroundocc; tpvformer). Driving scenes are highly sparse—most voxels are empty—so anisotropic Gaussians can shape and orient along semantic object surfaces, concentrating capacity where geometry and semantics lie while keeping free space compact, yielding an efficient 3D scene model. GaussianFormer (huang2024gaussianformer; huang2024gsformer2) represents scenes using sets of 3D Gaussians, each defined by a mean, covariance, and semantic label; these Gaussians are refined via deformable attention and rendered to a voxel grid through Gaussian-to-voxel splatting.
@@
In this work, we propose the first Vision-Only Gaussian Splatting framework for collaborative SOP that can be efficiently trained end-to-end in a single stage. Our method leverages sparse 3D semantic Gaussians as intermediate representations shared among vehicles, explicitly encoding both geometric and semantic information. By transmitting rigid transforms of Gaussians only within each ego vehicle’s region of interest (Figure 1 bottom), our approach effectively integrates multi-agent viewpoints, overcoming occlusions. Additionally, we introduce a novel cross-agent Gaussian fusion module that refines noisy, redundant, and conflicting Gaussians through neighborhood-based fusion, enabling accurate and realistic semantic occupancy prediction. To the best of our knowledge, we are the first to utilize 3D Gaussian splatting for multi-agent collaborative 3D semantic occupancy prediction. We summarize our contributions as:
@@
Semantic occupancy prediction (SOP), also known as semantic scene completion (SSC), estimates per-voxel occupancy together with semantic categories, yielding a volumetric representation of geometry and semantics. Since Tesla described an occupancy network for Full Self-Driving (tesla), research on 3D occupancy for autonomous driving has steadily accelerated. Methods differ by sensing modality: some use LiDAR point clouds (gaussianformer3d) or 4D radar tensors (4DRTs) (radarocc), while many recent systems are vision-only and lift image features into 3D using learned geometric priors (monoscene; ssc1; ssc2; ssc3; tpvformer; huang2024gaussianformer; surroundocc; occformer; voxformer). Most prior work addresses single-vehicle perception. Extending SOP to multi-agent collaborative perception centers on two design decisions: the communication representation and the fusion strategy for processing aggregated information under bandwidth limits and field-of-view differences. Collaborative SOP is only beginning to be explored; CoHFF (cohff) reports clear benefits from sharing but relies on multi-stage training and depth estimation, which increases system complexity and complicates cross-agent feature alignment.
@@
Collaborative perception for connected autonomous vehicles (CAVs) uses V2X communication and data fusion to improve scene understanding. Fusion is commonly grouped into early, intermediate, and late (cp_survey). Early fusion shares raw sensor data (e.g., point clouds), which is bandwidth demanding and raises privacy concerns (vieye). Late fusion exchanges object lists, which is lightweight but discards fine details (vips; otvic). Most recent systems adopt intermediate fusion, which exchanges latent features and retains more task signal while avoiding raw data transfer (v2vnet; where2comm; cobevt; v2xvit; v2xvitv2; cppc). While many works target LiDAR or hybrid setups for 3D detection or BEV segmentation (v2v4real; cppc; ermvp; tumtraf; disconet; where2comm; how2comm), vision-only collaboration is increasingly practical for large-scale deployment. CoHFF brought collaborative semantic occupancy prediction into focus by demonstrating SOP benefits from V2X feature sharing and fusion (cohff).
@@
3D Gaussian Splatting (3DGS) represents scenes with anisotropic Gaussians and renders them by splatting for novel view synthesis (3dgs). Subsequent work adapts Gaussian splatting to perception tasks, including object detection (3dgs-det; gaussiandet). For semantic occupancy, the GaussianFormer family splats Gaussians into voxels (huang2024gaussianformer; huang2024gsformer2). Prior work treats Gaussians as an internal single‑agent representation. We instead introduce Gaussians as the communication medium for collaboration. This choice provides compact messages, closed‑form rigid alignment across agents, region‑of‑interest culling, and explicit geometry that reduces reliance on separate depth supervision. Our method builds communication and cross-agent fusion for Gaussian primitives.
@@
$\mathbf{F}_{i}$ $=f_{\text{enc}}(\mathbf{O}_{i}),\kern 5.0pti\in\mathcal{S}\kern 5.0pt,$ (1)
$\mathbf{F}_{j\rightarrow i}$ $=\Gamma_{j\rightarrow i}(\mathbf{F}_{j}),\kern 5.0ptj\in\mathcal{S}\kern 5.0pt,$ (2)
$\mathbf{H}_{i}$ $=f_{\text{fusion}}(\{\mathbf{F}_{j\rightarrow i}\}_{j\in\mathcal{S}})\kern 5.0pt,$ (3)
$\mathbf{B}_{i}$ $=f_{\text{head}}(\mathbf{H}_{i})\kern 5.0pt,$ (4)
@@
$\max_{\theta}\sum_{i}g\left(\Phi_{\theta}(\mathbf{O}_{i},\{\mathbf{F}_{j\rightarrow i}\}_{j\in\mathcal{S}}),\mathbf{\hat{B}}_{i}\right),\kern 5.0pt\text{s.t.}\kern 5.0pt\sum_{j}\left|\mathbf{F}_{j\rightarrow i}\right|\leq\beta\kern 5.0pt,$ (5)
@@
Considering practical bandwidth limits, extending voxel-based single-agent SOP methods (e.g.,(surroundocc; voxformer)) to collaboration is not viable: transmitting dense voxel features in $\mathbb{R}^{X\times Y\times Z\times D}$ (where $D$ is the hidden dimension) is too costly. Drawn inspiration from TPVFormer (tpvformer), CoHFF proposes to transmit features from orthogonal planes $\mathbf{P}^{xz}$ and $\mathbf{P}^{yz}$, with features $\mathbf{F}^{\mathbf{P}^{xz}}\in\mathbb{R}^{X\times Z\times F}$ and $\mathbf{F}^{\mathbf{P}^{yz}}\in\mathbb{R}^{Y\times Z\times F}$. This reduces the communication volume from $XYZD$ to $(XZ+YZ)D$ and makes collaborative SOP feasible.
@@
Inspired by GaussianFormer (huang2024gaussianformer), we opt to represent a scene with a set of 3D Gaussian primitives $\mathcal{G}=\{\mathbf{G}_{i}\in\mathbb{R}^{d}|i=1,\ldots,P\}$. Distinct from GaussianFormer, we make the set of 3D Gaussian primitives serve as a communication medium for collaborative perception. Specifically, each $\mathbf{G}_{i}$ describes a local region with its mean $\mathbf{m}_{i}\in\mathbb{R}^{3}$, scale $\mathbf{s}_{i}\in\mathbb{R}^{3}$, rotation $\mathbf{r}_{i}\in\mathbb{R}^{4}$, opacity $a_{i}\in\mathbb{R}^{1}$, and semantics $\mathbf{c}_{i}\in\mathbb{R}^{|\mathcal{C}|}$. These primitives are interpreted as local semantic Gaussian distributions which contribute to the overall occupancy prediction for any given 3D location $\mathbf{x}\in\mathbb{R}^{3}$ through additive aggregation:
$\hat{\mathbf{o}}(\mathbf{x};\mathcal{G})=\sum_{i=1}^{P}\mathbf{g}_{i}(\mathbf{x};\mathbf{m}_{i},\mathbf{s}_{i},\mathbf{r}_{i},a_{i},\mathbf{c}_{i})\kern 5.0pt,$ (6)
@@
$\mathbf{g}(\mathbf{x};\mathbf{G})=a\cdot\exp\left(-\frac{1}{2}(\mathbf{x}-\mathbf{m})^{\text{T}}\mathbf{\Sigma}^{-1}(\mathbf{x}-\mathbf{m})\right)\mathbf{c}\kern 5.0pt,$ (7)
$\mathbf{\Sigma}=\mathbf{RSS}^{\top}\mathbf{R}^{\top}\kern 5.0pt,\kern 5.0pt\mathbf{S}=\text{diag}(\mathbf{s})\kern 5.0pt,\kern 5.0pt\mathbf{R}=\text{q2r}(\mathbf{r})\kern 5.0pt,$ (8)
@@
From a single-agent perspective, the pipeline first estimates 3D Gaussian primitives with an image-to-Gaussian module, then renders semantic occupancy via Gaussian-to-voxel splatting. We adopt both components from GaussianFormer (huang2024gaussianformer). We use 3D Gaussian primitives as a compact communication interface for collaborative perception. A Gaussian packaging module transforms each neighbor’s Gaussians into the ego frame and culls them to the ego region of interest, reducing the transmitted data. A cross-agent Gaussian fusion module then refines the ego set by aggregating consistent multi-view evidence and down-weighting noisy or low-opacity inputs. Collaborative semantic occupancy is obtained by splatting the refined ego Gaussians. The system is illustrated in Fig. 2.
Figure 2: Overview of the proposed pipeline. An initial set of randomly initialized 3D Gaussians is refined by an Image-to-Gaussian module (huang2024gaussianformer) that attends to multi-scale image features, producing single-agent Gaussians. Neighbor agents (top and bottom) are rigidly transformed into the ego frame (middle) and culled to the ego region of interest; the blue and yellow dashed box marks the Gaussians that lie within the ego ROI and are packaged and transmitted to the ego. A cross-agent Gaussian fusion module aggregates these with the ego set. The fused Gaussians are then rendered to semantic occupancy via Gaussian-to-voxel splatting (huang2024gaussianformer). For clarity, the figure shows the zero-shot variant.
@@
Unlike dense voxel grid features (surroundocc; voxformer) or planar features (cohff; tpvformer), Gaussian primitives model ellipsoidal probability densities in $\mathbb{R}^{3}$, which remain well defined under rigid transforms (3dgs; huang2024gaussianformer), and thus cross-agent 3D Gaussian alignment and transmission are straightforward and interpretable. Specifically, let the known extrinsic from agent $j$ to agent $i$ be $T_{ij}(x)=U_{ij}x+t_{ij}$ with $U_{ij}\in\mathrm{SO}(3)$, $t_{ij}\in\mathbb{R}^{3}$, and quaternion $q_{ij}$ such that $\text{q2r}(q_{ij})=U_{ij}$.
@@
$\mathbf{m}^{\prime}$ $=U_{ij}\,\mathbf{m}+t_{ij}\kern 5.0pt,$ (9)
$\mathbf{\Sigma}^{\prime}$ $=U_{ij}\,\mathbf{\Sigma}\,U_{ij}^{\top}\;=\;\big(U_{ij}\mathbf{R}\big)\,\mathbf{S}\mathbf{S}^{\top}\,\big(U_{ij}\mathbf{R}\big)^{\top}\kern 5.0pt.$ (10)
@@
$\mathbf{m}^{\prime}=U_{ij}\mathbf{m}+t_{ij}\kern 5.0pt,\kern 5.0pt\mathbf{s}^{\prime}=\mathbf{s}\kern 5.0pt,\kern 5.0pt$
$\mathbf{r}^{\prime}=q_{ij}\otimes\mathbf{r}\kern 5.0pt,\kern 5.0pta^{\prime}=a\kern 5.0pt,\kern 5.0pt\mathbf{c}^{\prime}=\mathbf{c}\kern 5.0pt,$ (11)
@@
$\mathcal{G}_{j\rightarrow i}\;=\;\{\,\mathbf{G}\in\mathcal{G}_{j}\;|\;\mathbf{m}^{\prime}\in\mathbf{ROI}_{i}\,\}\kern 5.0pt.$ (12)
@@
$\mathcal{G}^{\mathrm{stack}}_{i}\;=\;\mathcal{G}_{i}\;\cup\;\{\mathcal{G}_{j\rightarrow i}\}_{j\in\mathcal{S}}\kern 5.0pt,$ (13)
@@
Unlike the refinement step in the image-to-Gaussian module of GaussianFormer (huang2024gaussianformer), our fusion module decodes neighborhood-conditioned proposals, pools them across nearby Gaussians to suppress noise and enforce consistency, and then blends the pooled update with the ego Gaussian attributes.
@@
Table 1: Comparison for semantic occupancy prediction. In the collaborative setting, our naive fusion and learned fusion variants achieve better IoU and mIoU than CoHFF (cohff). The best is in bold and the second best is underlined.
@@
$\mathcal{G}_{\text{nbgr},k}=\{\,\mathbf{G}_{j}\in\mathcal{G}_{\text{nbgr}}\mid\|\mathbf{m}_{j}-\mathbf{m}_{k}\|_{2}\leq\rho\,\}\kern 5.0pt.$ (14)
@@
$\mathbf{f}^{\mathrm{ego}}_{k}$ $=\big[\,\mathbf{m}_{k}^{\top},\ \mathbf{s}_{k}^{\top},\ \mathbf{r}_{k}^{\top},\ a_{k},\ \mathbf{c}_{k}^{\top}\,\big]\kern 5.0pt,$ (15)
$\mathbf{f}^{\mathrm{rel}}_{k,j}$ $=\big[\,(\mathbf{m}_{j}{-}\mathbf{m}_{k})^{\top},\ (\mathbf{s}_{j}{-}\mathbf{s}_{k})^{\top},\ |\langle\mathbf{r}_{j},\mathbf{r}_{k}\rangle|,\ a_{j},\ \mathbf{c}_{j}^{\top}\,\big]\kern 5.0pt,$
$\mathbf{z}_{k,j}$ $=\big[\,\mathbf{f}^{\mathrm{ego}}_{k}\ \big|\ \mathbf{f}^{\mathrm{rel}}_{k,j}\,\big]\kern 5.0pt,$
@@
$\mathbf{u}_{k,j}\coloneq\big(\Delta\mathbf{m}_{k,j},\ \mathbf{s}^{\star}_{k,j},\ \mathbf{r}^{\star}_{k,j},\ a^{\star}_{k,j},\ \mathbf{c}^{\star}_{k,j}\big)\;=\;\text{MLP}(\mathbf{z}_{k,j})\kern 5.0pt,$ (16)
@@
$w_{k,j}=\mathrm{softmax}_{j}\!\left(\frac{\langle Q\,\mathbf{x}^{\text{ego}}_{k},\ K\,\mathbf{x}^{\text{rel}}_{k,j}\rangle}{\sqrt{d}}\right)\kern 5.0pt,$ (17)
@@
$\bar{\mathbf{u}}_{k}=\sum_{G_{j}\in\mathcal{G}_{\text{nbgr},k}}w_{k,j}\,\mathbf{u}_{k,j}\kern 5.0pt.$ (18)
@@
$\hat{\mathbf{m}_{k}}$ $=\mathbf{m}_{k}+\overline{\Delta\mathbf{m}}_{k}\kern 5.0pt,\qquad\hat{\mathbf{s}_{k}}\;=\;\overline{\mathbf{s}^{\star}}_{k}\kern 5.0pt,$ (19)
$\hat{\mathbf{r}_{k}}$ $=\overline{\mathbf{r}^{\star}}_{k}\kern 5.0pt,\qquad\hat{a}_{k}\;=\;\overline{a^{\star}}_{k}\kern 5.0pt,$
$\hat{\mathbf{c}_{k}}$ $=\alpha_{k}\,\mathbf{c}_{k}+(1{-}\alpha_{k})\,\overline{\mathbf{c}^{\star}}_{k}\kern 5.0pt,$
@@
$\alpha_{k}=\frac{\mathrm{conf}(\mathbf{c}_{k})}{\mathrm{conf}(\mathbf{c}_{k})+\mathrm{conf}(\overline{\mathbf{c}^{\star}}_{k})}\,\kern 5.0pt,\quad\mathrm{conf}(v)=\max\!\Big(\frac{v}{\mathbf{1}^{\top}v+\varepsilon}\Big)\kern 5.0pt.$ (20)